OpenStudy (nthenic_oftime):

Use the definition of continuity to determine whether f is continuous at a. f(x) = 5x4 - 9x3 + x - 7a = 7

OpenStudy (anonymous):

what grade is this anyway

OpenStudy (nthenic_oftime):

its a pre cal class

OpenStudy (nthenic_oftime):

can you help?

OpenStudy (jango_in_dtown):

f(x)=5x^4-9x^3+x-7a-7??

OpenStudy (anonymous):

f(0) = 0 - 4 = -4 lim (x->0, x>0) f(x) = 0^2 + 0 - 4 = -4 f is continuous.

OpenStudy (jango_in_dtown):

@nthenic_oftime ??

OpenStudy (jango_in_dtown):

@Roseamora it is told to find continuity at x=a

OpenStudy (nthenic_oftime):

thats what the proble said

OpenStudy (nthenic_oftime):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

see every polynomial function is continuous. can you give me a snapshot of the problem from the book?

OpenStudy (nthenic_oftime):

Use the definition of continuity to determine whether f is continuous at a. f(x) = 5x4 - 9x3 + x - 7a = 7 A. Not continuous B. Continuous @jango_IN_DTOWN this is the whole problem i am confused

OpenStudy (jango_in_dtown):

f(x)=5x^4-9x^3+x-7a-7 f(a)= 5a^4-9a^3+a-7a-7=5a^4-9a^3-6a-7= lim x->a f(x) hence continuous